3 July 2017
Monday 3rd of July 2017

Jan, Ina and Wataru are the happy physical properties team. They have just finished the overview measurements of the upper 350 meters of EGRIP ice core.

A busy day at EGRIP

Today another milestone was reached here at EGRIP: The physical properties team completed the overview measurements of c-axis and grain size for the upper 350 meters of the EGRIP core. As the main purpose of the EGRIP is to understand the NEGIS ice stream and its dynamics, measurements of the physical properties of the ice are absolutely essential.

In the drill trench we had to stop the drilling in the morning as the electronics section had malfunctioned and thus needed the attention of Bent once more. Absent new ice core, we decided to use the day for catching up on a number of big and small issues that have accumulated during the last days, making for a busy day here at EGRIP. In the logging cabin, heater were brought in, to help with the defrosting of the cooling unit, resulting in removal of approx. 95% of the ice that had accumulated in the unit, also the vacuum system for cleaning the cores was serviced. Between the core buffer and the science trench, new heavier curtains were installed to maintain a core buffer temperature below -20C. The loggers also helped out in cleaning core troughs to accelerate that process. The drillers made a new bore hole logging, to have a better understanding of the logging uncertainties, and they also made improvements in the drilling software to get more accurate numbers for inclination during drill runs and to put in more guidance for the drilling operation. Towards the evening the repaired electronics section was tested and re-installed in the drill.

The surface program and the water isotope measurements also made steady progress during the day.

What we did today:

  1. Servicing snowmobiles, moving drill liquid to the trench.
  2. Celebrated Vasileios’s birthday (Instruments: Flute, Tuba and Drums).
  3. Deep drilling core length Monday at 19.00: 7.33 m.
  4. Logging depth at 19.00, Sunday: 573.12 m.
  5. Bore hole logging.
  6. Repair of new electronics section, improvements implemented in drilling software.
  7. Heating logging cabin so the logging cooler could be almost entirely defrosted.
  8. Water isotopes: Melting of bags 40 to 59.
  9. Water vapour, methane, aerosols, snow and boundary layer measurements
  10. Physical properties measurements, finished up overview sampling of c-axis and grain sizes.
  11. Installed new (heavy) curtains between core buffer and science trench
  12. Cleaning of core troughs.

Weather: The day started with some ground fog, turning overcast and ending with sunny spells. Light winds from Easterly directions throughout the day. Temperature between -18°C and -13°C.

FL, Bo Vinther

← Previous entry   Next entry →